OpenStudy (noseboy908):

Solve along the interval [theta between 0 and 2pi]" sin(3theta)=-1

OpenStudy (3mar):

do you mean \[\sin(3 \theta)=-1\]?

OpenStudy (noseboy908):

A little more eloquently put:\[\sin (3\theta)=-1\]

OpenStudy (3mar):

OK. Let's draw  What do you think the correct angle matches this equation?

OpenStudy (noseboy908):

The bottom, where the ordered pair is (0,-1).

OpenStudy (3mar):

Great!

OpenStudy (3mar):

so we can say that: \[3\theta=\frac{ -\pi }{ 2 } +n.2\pi\] if theta in radian.

OpenStudy (3mar):

and \[3\theta=-90+n*360\] if it is in degree.

OpenStudy (3mar):

Do you follow?

OpenStudy (noseboy908):

Explain how you get "+n.2pi" above, and where n came from.

OpenStudy (3mar):

Firstly, your theta in radian or degree?

OpenStudy (noseboy908):

Radians.

OpenStudy (3mar):

Ok Origanally, your answer would be \[3\theta=\frac{ -\pi }{ 2 }\] \[\theta=\frac{ -\pi }{ 6 }\] and we know that every angle theta is equal to its sisters with the addition of 2pi and 4pi and 6pi ..... e.g: if a=30 then a=30+2pi=30+4pi=30+6pi=30+n(2pi) where n is 0,1,2,3,.....

OpenStudy (noseboy908):

Ok, that makes sense

OpenStudy (3mar):

Did you get it?

OpenStudy (noseboy908):

I understand all you said there, but you made no mention of how you incorporated the "n' in it.

OpenStudy (3mar):

Yes, I did. "if a=30 then a=30+2pi=30+4pi=30+6pi=30+n(2pi) where n is 0,1,2,3,....."

OpenStudy (noseboy908):

I understand

OpenStudy (3mar):

So can you share me the final result you are comfortable with?

OpenStudy (noseboy908):

Is the the same as saying\[\sin \theta=-1/3\] ?

OpenStudy (3mar):

Not quite right. \[\sin(3\theta)=-1\] \[\sin^{-1}(-1)=3\theta\] \[3\theta=\frac{ -\pi }{ 2 }\] \[\theta=\frac{ -\pi }{ 6 }\] and if you want: \[\theta=\frac{ -\pi }{ 6 }+n.\frac{ 2\pi }{ 3 }\] but don't care of that if it would confuse you!

OpenStudy (noseboy908):

Got it

OpenStudy (3mar):

Thank God!

OpenStudy (3mar):

Thank you for the medal!

OpenStudy (noseboy908):

My feeling as well. Enough math for one day.

OpenStudy (phi):

notice they want only the solutions the interval [theta between 0 and 2pi]" your solution is for all values of theta, so you have to pick out only the few between 0 and 2pi